# LLaVA-UHD v2: an MLLM Integrating High-Resolution Feature Pyramid via Hierarchical Window Transformer

### v1

@@ Abstract
In multimodal large language models (MLLMs), vision transformers (ViTs) are widely employed for visual encoding. However, their performance in solving universal MLLM tasks is not satisfactory. We attribute it to a lack of information from diverse visual levels, impeding alignment with the various semantic granularity required for language generation. To address this issue, we present LLaVA-UHD v2, an advanced MLLM centered around a Hierarchical window transformer that enables capturing diverse visual granularity by constructing and integrating a high-resolution feature pyramid. As a vision-language projector, Hiwin transformer comprises two primary modules: (i) an inverse feature pyramid, constructed by a ViT-derived feature up-sampling process utilizing high-frequency details from an image pyramid, and (ii) hierarchical window attention, focusing on a set of key sampling features within cross-scale windows to condense multi-level feature maps. Extensive experiments demonstrate that LLaVA-UHD v2 achieves superior performance over existing MLLMs on popular benchmarks. Notably, our design brings an average boost of 3.7% across 14 benchmarks compared with the baseline method, 9.3% on DocVQA for instance. We make all the data, model checkpoint, and code publicly available to facilitate future research.
@@ 1 Introduction
Figure 1: Comparison of LLaVA-UHD v2 with other MLLMs. (a) MLLMs typically align ViT features to language space using MLPs [63] or perceiver re-samplers [6, 52], lacking visual granularity. (b) Combining multiple visual encoders is non-universal and computationally intensive. (c) LLaVA-UHD v2 employs the Hiwin transformer to build an inverse feature pyramid and compress it into visual tokens, providing various semantic granularity for language generation.
@@
Embedding visual information into large language models (LLMs) [97, 21, 4, 2] has significantly enhanced their ability to tackle complex visual multimodal tasks, such as visual question answering [8, 38, 32], document analysis [78, 77, 66], and visual interaction [19, 27]. With the rise of transformers [98], the vision transformer (ViT) [24, 85] has emerged as a standard approach for visual encoding in contemporary multimodal large language models (MLLMs) [23, 10, 99, 63, 62, 52, 4, 7, 103].
However, the ViT in an MLLM struggles with effectively handling all types of visual multimodal tasks, due to inadequate visual granularity captured by single-scale features, as illustrated in Fig. 1(a). For instance, visual grounding [107, 17, 110, 106] and optical character recognition [46, 49] demand fine-grained visual details, while image caption [84, 57] focuses on capturing high-level semantics and relations.
Drawing on the advancements in computer vision, comprehensive coverage of fine-grained details and high-level semantics can be integrated effectively by a feature pyramid [59, 87, 114]. However, this method presents two key challenges when applied to MLLMs: (1) Representation Inheritance. With its scalability, the ViT boasts general vision-language representations acquired through large-scale image-text pre-training ($e.g.$, CLIP [85] or SigLIP [109]), which is essential for continual learning in MLLMs. Modifying the ViT into a hierarchical architecture such as Swin [67] for multi-scale representations renders the vision-language representations of existing pre-trained models non-heritable. (2) Compression Effectiveness. The quadratic computational cost of LLMs with respect to the number of visual tokens necessitates effective compression of the feature pyramid. Current techniques [91, 60, 71, 96] often resize features to a fixed resolution, which distorts both the spatial shape and intrinsic semantics, as shown in Fig. 1(b). Moreover, down-sampling high-resolution features risks losing critical fine-grained details.
@@
To address these challenges, we present LLaVA-UHD v2, an advanced MLLM centered around a novel Hierarchical window transformer, which enables capturing diverse visual granularities by constructing and integrating a high-resolution feature pyramid, as shown in Fig. 1(c). The Hiwin transformer performs an expansion-then-compression procedure on spatial resolution of feature maps, which consists of two stages: (i) Constructing an inverse feature pyramid. We train a parameterized feature up-sampler, the Joint Bilateral Upsampling (JBU) module proposed in FeatUp [26], on top of a pre-trained ViT ($i.e.$, CLIP-ViT [85]) to encode high-frequency information from the image into its feature map. This strategy can be applied to any ViT, inheriting its powerful feature representations, while constructing a multi-resolution feature pyramid based on the ViT. (ii) Integrating features with hierarchical window attention. To condense the multi-level feature maps effectively, we propose utilizing a set of hierarchical windows to capture semantics from local regions across different pyramid levels. A set of learnable queries is restricted to attend solely to key sampled features within their respective windows. This attention mechanism performs effective compression on local dense features at the native resolution of each pyramid level, thereby enabling visual tokens to effectively capture both fine-grained details and high-level semantics.
@@
Extensive experiments demonstrate that LLaVA-UHD v2 dramatically outperforms the baseline method (LLaVA-UHD [31]) across 14 popular benchmarks, including document-centric visual question answering ($e.g.$, +9.3% on DocVQA), visual grounding ($e.g.$, average +5.7% on RefCOCOs [107]), and high-resolution image perception ($e.g.$, +3.4% on HR-Bench [100]). Moreover, we experimentally reveal that feature pyramids, regardless of the specific construction method ($e.g.$, bilinear interpolation), can further enhance the visual perception capabilities of MLLMs, offering new insights for future research. Our contributions can be concluded as follows:
  • We propose a visual pyramid encoding method by pre-training an up-sampling module, providing enriched semantic granularity for language decoding.
  • We propose a hierarchical window attention, capable of integrating and compressing a high-resolution feature pyramid into a condensed set of visual tokens, thereby ensuring computational efficiency while preserving the critical details needed for language decoding.
@@
  • Trained efficiently on academic-scale data, LLaVA-UHD v2 achieves substantial improvements over the baseline method across 14 benchmarks.
@@ 2 Related Works
Figure 2: The overall architecture of proposed LLaVA-UHD v2, consisting of a ViT, our hierarchical window transformer (Hiwin transformer), and an LLM. Hiwin transformers process sliced patches and the overview image by capturing inner multi-level representations and compressing them into spatially consistent tokens for a better vision-language alignment.
@@ 2.1 Feature Pyramid Representation
Image pyramid techniques are fundamental in image processing, facilitating multi-resolution analysis since the era of manual feature design, as seen in SIFT’s scale-space keypoints [13, 70]. In deep learning, CNNs such as ResNet and VGG, inherently extract hierarchical features across scales [35, 93]. Innovations such as FPN and U-Net enhance semantic hierarchy for tasks like detection and segmentation [59, 87]. Recently, some transformer-based models [67, 111, 114] further advanced feature pyramid construction, capturing more comprehensive visual semantic granularity for visual representations. However, multimodal language models, often using CLIP-based ViTs, underutilize hierarchical features, suggesting a research gap for integrating advanced feature pyramids into these models [111, 85].
@@ 2.2 Visual Encoding in MLLMs
CLIP-ViT, favored for its effective alignment of visual features with linguistic semantics through contrastive pre-training, is widely adopted in MLLMs [85, 64, 63, 10, 62, 7, 113, 55]. Emerging research explores alternative visual representations, primarily in three categories: (1) Fusing features from CLIP-based CNNs and ViTs. LLaVA-HR [73] integrates stage-wise CNN features [68] into ViT’s layers, enhancing fine-grained perception. CogAgent [37] utilizes ViT features to query high-resolution CNN features for detailed information during language decoding. Mini-Gemini [54] employs a cross-attention-based post-fusion between CNN and ViT features. (2) Fusing features from visual experts trained with different vision pre-training tasks [60, 28, 71, 115, 96, 91, 102]. Candidate experts include DINO-v2 [83] by visual contrastive pre-training, SAM [47] by prompt segmentation pre-training, Pix2Struct [49] by document parsing pre-training, $etc.$ Deepseek-VL [71], SPHINX-X [60, 28] and Eagle [91] down-sample output feature maps and concatenate them along the channel axis. Cambrian-1 [96] initializes embeddings to query local patches from visual experts. (3) Language models for visual encoding: Fuyu [11], Otter-HD [51], and SOLO [20] encode images directly with LLMs, bypassing dedicated visual encoders. However, these approaches, while effective, increase computational demand and hinder a unified image-to-language design.
@@ 3.1 Overview
The architecture of the proposed LLaVA-UHD v2 is illustrated in Fig. 2. It consists of three primary modules: a visual encoder (ViT), a vision-language projector (Hiwin transformer), and an LLM. We first divide and crop each input image into multiple slices with appropriate size and aspect ratio, leveraging the slicing strategy proposed in [31]. The image slices, along with an overview image, are then processed by a CLIP-ViT [85] with interpolated positional embeddings, capable of processing images with arbitrary sizes and shapes. The resulting features are subsequently passed to the Hiwin transformer for vision-language projection, which is carried out with two stages: (i) constructing an inverse feature pyramid (detailed in Sec.3.2), and (ii) integrating the feature pyramid by hierarchical window attention (detailed in Sec.3.3). The core of the Hiwin transformer lies in enhancing each ViT-encoded feature into a high-resolution feature pyramid encoding, thereby achieving enriched semantic granularity for each image slice. After this enhancement, visual tokens from different slices are reorganized into a spatially consistent feature map relative to the original image, ensuring clarity in spatial relationships. Following by concatenating with the overview tokens, all the visual tokens provide visual context for language decoding.
3.2 Inverse Feature Pyramid
@@
Preliminaries. In the era of hand-crafted visual representations [70, 13], gaussian convolution filter is widely employed to construct image pyramids $\{\mathcal{I}^{l}\in\mathbb{R}^{\frac{H}{2^{l}}\times\frac{W}{2^{l}}\times 3},l=0,1...L-1\}$, where $l$ denotes the pyramid level, $(H,W)$ the image height and width, and $L$ the total number of levels. Building upon this concept, deep convolutional neural networks (CNNs), such as ResNet and VGG, naturally yield a bottom-up, hierarchical feature pyramid $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H}{p\cdot 2^{l}}\times\frac{W}{p\cdot 2^{l}}\times C}\}$, where $C$ represents the feature dimension and $p$ denotes the a down-sampling ratio ($e.g.$, $p=8$ in ResNet-50 [35]). In such pyramids, lower-level feature maps generally have higher resolution and capture more visual detail, whereas higher-level feature maps exhibit lower resolution but contain more abstract semantic information. However, ViTs splitting the image into coarse patches, only produce single-scale feature maps ($i.e.$, $\mathcal{F}^{0}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times C}$, where for instance, $p=14$), as illustrated in Fig. 1(a). Lacking such feature pyramid, like that present in CNNs, hinders the performance of ViTs on MLLM tasks, which require both fine-grained visual details and high-level semantic information. Consequently, how to construct a ViT-based feature pyramid with varying semantic granularity remains an unresolved challenge.
Figure 3: Flowchart of the Joint Bilateral Upsampling (JBU) module, which leverages the image pyramid to guide feature up-sampling, integrating high-frequency information into the up-sampled feature maps.
@@
Guided feature up-sampling. Due to the lack of high-resolution features, up-sampling ViT features becomes the necessary strategy to inversely construct the feature pyramid. Two straightforward approaches can be employed:(1) plain bilinear interpolation and (2) deconvolution network. By doubling and quadrupling the resolution of resulting feature maps, a ViT-based feature pyramid $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times C},l=0,1,2\}$ is constructed. While effective, directly up-sampling deep features hardly introduces precise visual details, which are essential for tasks requiring fine-grained visual information. To address this limitation, we leverage the parameterized JBU module [26] to guide feature up-sampling with image priors, which helps introduce high-frequency visual details.
@@
Generally, the objective of JBU module is to learn $(l-1)$ convolutional layers on the image pyramid $\{\mathcal{I}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times 3}\}$ to capture high-frequency patterns of local texture for guiding feature up-sampling, in Fig. 3. Specifically, for each input image, its $(l+1)$-th level ViT features is defined as
@@
where $\mathrm{Up}(\cdot)$ denotes the up-sampling interpolation and $\mathrm{Conv}(\cdot)$ represents the convolutional operation applied to feature maps using customized kernel weights $\Theta^{l+1}$ learned from the corresponding image $\mathcal{I}^{l+1}$.
As a result, each pixel value of the up-sampled feature maps is extracted as
where $[x,y]$ denotes the coordinate index of feature maps, $U$ the region of convolution neighborhood, $D_{dist}$ is a decay factor based on the distance between $[x,y]\ \text{and}\ [x^{\prime},y^{\prime}]$, and $D_{sim}$ is a similarity weight based on the attention of each image pixel using a two-layer GeLU MLP $\Theta^{l+1}$. Additional details can be found in supplement A.1. By iteratively performing Eq. 1, we progressively up-sample the feature maps and ultimately construct the feature pyramid $\{{\mathcal{F}^{l},l=0,1,2}\}$.
Learnable up-sampler optimizing. Following [26], we optimize the reconstruction loss between the highest-level feature maps and the lowest as
$\mathcal{L}=\lVert\mathcal{F}^{0}-\mathrm{Down}(\mathcal{F}^{2};\Omega)\rVert_{2}^{2},$ (3)
where $\mathrm{Down}(\cdot)$ is a down-sampling network with trainable weights $\Omega$, $e.g.$, two convolution layers with a stride of 2. Note that we omit some hyper-parameters in Eq. 3 for clarity. Please kindly refer to the supplement A.1 for details.
However, when training the JBU module upon CLIP-ViT [85], we encounter the texture detail degradation of high-resolution feature maps. Specifically, features of CLIP-ViT scarcely exhibit enhanced detail representation at higher resolutions, compared to other ViTs like DINO-ViT [83]. We attribute this issue to the lack of inter-level feature reconstruction for language-aligned features during iterative up-sampling, resulting in an unstable transformation of such semantic-centric representations of the CLIP-ViT. To address this, we propose a straightforward hierarchical supervision strategy, regulating each level in the pyramid. Eq. 3 is updated as follows. The complete form for Eq. 3 and additional details can be found in supplement A.1.
$\mathcal{L}=\frac{1}{2}\sum_{l=1}^{2}\lVert\mathcal{F}^{0}-\mathrm{Down}(\mathcal{F}^{l};\Omega^{l})\rVert_{2}^{2}$ (4)
Figure 4: The flowchart of hierarchical window attention. Feature maps from different levels of the feature pyramid are adaptively RoI-aligned into sampling features and then concatenated along the length axis to serve as the key for the learnable queries.

### v2

# LLaVA-UHD v2: an MLLM Integrating High-Resolution Semantic Pyramid via Hierarchical Window Transformer

@@ Abstract
Vision transformers (ViTs) are widely employed in multimodal large language models (MLLMs) for visual encoding. However, they exhibit inferior performance on tasks regarding fine-grained visual perception. We attribute this to the limitations of ViTs in capturing diverse multi-modal visual levels, such as low-level details. To address this issue, we present LLaVA-UHD v2, an MLLM with advanced perception abilities by introducing a well-designed vision-language projector, the Hierarchical window (Hiwin) transformer. Hiwin transformer enhances MLLM’s ability to capture diverse multi-modal visual granularities, by incorporating our constructed high-resolution semantic pyramid. Specifically, Hiwin transformer comprises two key modules: (i) a visual detail injection module, which progressively injects low-level visual details into high-level language-aligned semantics features, thereby forming an inverse semantic pyramid (ISP), and (ii) a hierarchical window attention module, which leverages cross-scale windows to condense multi-level semantics from the ISP. Extensive experiments show that LLaVA-UHD v2 outperforms compared MLLMs on a wide range of benchmarks. Notably, our design achieves an average boost of 3.7% across 14 benchmarks compared with the baseline method, 9.3% on DocVQA for instance. All the data and code will be publicly available to facilitate future research.
@@ 1 Introduction
Figure 1: Comparison of LLaVA-UHD v2 with other MLLMs. (a) MLLMs typically align ViT features to language space using MLPs [67] or perceiver re-samplers [6, 56], lacking visual granularity. (b) Combining multiple visual encoders is non-universal and computationally intensive. (c) LLaVA-UHD v2 employs the Hiwin transformer to build an inverse semantic pyramid and compress it into visual tokens, providing various semantic granularity for language generation.
@@
Embedding visual information into large language models (LLMs) [103, 23, 4, 2] has significantly enhanced their ability to handle multimodal tasks, such as visual question answering [9, 42, 36], document analysis [85, 84, 70], and visual interaction [21, 30].
Among these advancements, the CLIP [91] series, built upon the vision transformer (ViT) architecture [27, 91], has emerged as a standard for visual encoding in contemporary multi-modal large language models (MLLMs) [26, 11, 104, 67, 66, 56, 4, 8, 110]. By leveraging contrastive training, CLIP-ViT extracts text-aligned visual features that facilitate seamless integration with LLMs, enabling effective handling of tasks that emphasize text-aligned semantics, $e.g.$, image captioning [90, 61].
However, CLIP-ViT-based MLLMs often underperform in tasks requiring extensive low-level visual details, such as visual grounding [114, 18, 116, 113] and optical character recognition [50, 53], hindering the MLLMs’ practical applications. As illustrated in Fig. 1(a), such inferior performance primarily stems from the low resolution and text-aligned nature of CLIP features. To facilitate MLLMs’ fine-grained perception, Yao et al. [109] proposes a Dense Connector [109], which fuse lower layers of CLIP-ViT. Yet, its performance is limited by the low resolution of the features.
To address such limitation, we attempts to incorporate a vision feature pyramid [63, 93, 122] into MLLMs to provide multi-level visual granularity. However, directly employing such pyramid might encounter the following two key challenges: (1) Representation Non-inheritance. The inherent image-text alignment of CLIP-ViT’s feature representations is fundamental to the efficacy of MLLMs. Substituting CLIP-ViT with hierarchical architectures such as Swin [72] to obtain multi-scale representations would disrupt this alignment and consequently compromise the transferability of its pretrained representations. (2) Compression Ineffectiveness. The quadratic computational cost of LLMs w.r.t. the number of visual tokens necessitates effective compression of the feature pyramid. Current projecting methods [97, 64, 77, 102] designed a customized projector to resize multi-scale visual features before token compression, which resulted in a loss in visual details and spatial relations in naive feature resolution, as shown in Fig. 1(b).
To address these issues, we present LLaVA-UHD v2, an advanced MLLM with enhanced perceptual capabilities leveraging the incorporated Hiwin transformer. As shown in Fig. 1(c), Hiwin transformer enables capturing diverse multi-modal granularity by integrating a constructed high-resolution semantic pyramid.
@@
Specifically, the Hiwin transformer consists of two key modules as follows: (i) a visual detail injection module (VDIM) for inheriting the representations. We propose a VDIM to inject low-level details ($e.g.$, edges, textures) from images into text-aligned features from a CLIP-pretrained ViT [91], progressively building an up-sampled level upon the previous level, resulting in an inverse semantic pyramid (ISP). During the training stage, a reconstruction loss between fused and original CLIP features explicitly maintains vision-language alignment while enhancing visual granularity. This strategy can be extended to any ViT for inheriting its powerful multi-modal representations. (ii) a hierarchical window attention module for ensuring effective compression. We propose utilizing a set of hierarchical windows to capture semantics from local regions across different semantic levels. A set of learnable queries is restricted to attending solely to sampled features within their respective windows. This attention mechanism performs effective compression on local dense features at the native resolution of each pyramid level, thereby enabling visual tokens to effectively capture both fine-grained visual details and high-level language-aligned semantics.
@@
Extensive experiments demonstrate that LLaVA-UHD v2 dramatically outperforms the compared MLLMs across a wide range of benchmarks. More importantly, LLaVA-UHD v2 surpasses the baseline method (LLaVA-UHD [35]) on 14 popular benchmarks by 3.7% in average, including document-centric visual question answering ($e.g.$, +9.3% on DocVQA), visual grounding ($e.g.$, average +5.7% on RefCOCOs [114]), and high-resolution image perception ($e.g.$, +3.4% on HR-Bench [105]). Besides, we experimentally reveal that ISP, regardless of construction method ($e.g.$, bilinear interpolation), enhances the visual perception capabilities of MLLMs, offering new insights for future research. In summary, our contributions are three-fold:
  • We present LLaVA-UHD v2, an MLLM with the advanced visual perception ability, through the integration of a high-resolution semantic pyramid representation.
  • We propose the Hiwin transformer, a novel vision-language projector comprises a visual detail injection module and a hierarchical window attention module for capturing diverse multimodal visual granularities.
@@
  • Trained on merely academic-scale data, LLaVA-UHD v2 achieves substantial improvements over the baseline method across 14 benchmarks.
@@ 2 Related Works
Figure 2: The overall architecture of proposed LLaVA-UHD v2, consisting of a ViT, our hierarchical window transformer (Hiwin transformer), and an LLM. The Hiwin transformer first injects high-frequency visual details from the image into the high-level semantics of ViT features, forming inverse semantic pyramids (ISP). Then it compresses the ISPs into spatially consistent tokens via cross-scale windows, for a better vision-language alignment. Details about the two procedures are illustrated in Figure 3 and 4.
@@ 2.1 Feature Pyramid Representation
Image pyramid techniques are fundamental in image processing, facilitating multi-resolution analysis since the era of manual feature design, as seen in SIFT’s scale-space keypoints [14, 75]. In deep learning, CNNs such as ResNet and VGG, inherently extracts hierarchical features across scales [39, 99]. Innovations such as FPN and U-Net enhance semantic hierarchy for tasks like detection and segmentation [63, 93]. Recently, some transformer-based models [72, 117, 122] further advanced feature pyramid construction, capturing more comprehensive visual semantic granularity for visual representations. However, multimodal language models, often using CLIP-based ViTs, underutilize hierarchical features, suggesting a research gap for integrating advanced feature pyramids into these models [117, 91].
@@ 2.2 Visual Encoding in MLLMs
CLIP-ViT, favored for its effective alignment of visual features with linguistic semantics through contrastive pre-training, is widely adopted in MLLMs [91, 68, 67, 11, 66, 8, 121, 59]. Emerging research explores alternative visual representations, primarily in three categories: (1) Fusing features from CLIP-based CNNs and ViTs. LLaVA-HR [80] integrates stage-wise CNN features [73] into ViT’s layers, enhancing the fine-grained perception of ViT representations. CogAgent [41] utilizes ViT features to query high-resolution CNN features for detailed information during language decoding. Mini-Gemini [58] employs a cross-attention-based post-fusion between CNN and ViT features. (2) Fusing features from visual experts trained with different vision pre-training tasks [64, 31, 77, 123, 102, 97, 108]. Candidate experts include DINO-v2 [89] by visual contrastive pre-training, SAM [51] by prompt segmentation pre-training, Pix2Struct [53] by document parsing pre-training, $etc.$ Deepseek-VL [77], SPHINX-X [64, 31] and Eagle [97] down-sample output feature maps and concatenate them along the channel axis. Cambrian-1 [102] initializes embeddings to query local patches from visual experts. (3) Language models for visual encoding: Fuyu [12], Otter-HD [55], and SOLO [22] encode images directly with LLMs, bypassing dedicated visual encoders. However, these approaches, while effective, increase computational demand and hinder a unified image-to-language design.
@@ 3.1 Overview
The architecture of the proposed LLaVA-UHD v2 is illustrated in Fig. 2. LLaVA-UHD v2 comprises three modules: a visual encoder (ViT), a vision-language projector (Hiwin transformer), and an LLM. We first partition and process the input image with the adaptive slicing strategy from [35], outputting CLIP [91] features with arbitrary size and shape. The resulting features are subsequently passed to the Hiwin transformer for vision-language projection, which is carried out with two stages: (i) constructing an inverse semantic pyramid (ISP) and (ii) integrating the ISP by hierarchical window attention, which will be detailed in Sec.3.2 and Sec.3.3. The core of the Hiwin transformer lies in enhancing each ViT-encoded feature into a high-resolution semantic pyramid encoding, thereby achieving enriched semantic granularity for each slice or image. After the enhancement, visual tokens from different slices are reorganized into a spatially consistent feature map relative to the original image, ensuring clarity in spatial relationships. Followed by concatenating with the overview tokens, the visual tokens provide both high-level language-aligned semantics and high-resolution visual details for language decoding.
Figure 3: The flowchart illustrates the construction of the Inverse Semantic Pyramid (ISP). As the first level of ISP, $\mathcal{F}^{0}$ is the high-level language-aligned semantic features from CLIP-ViT. Subsequent levels, $\mathcal{F}^{1}$ and $\mathcal{F}^{2}$, are progressively built by injecting high-frequency visual details from the input image into upsampled features from the previous level, via the Visual Detail Injection Module (VDIM). A Multi-level Reconstruction (MLR) loss supervises in each scale, ensuring both text-aligned semantic coherence and fine-grained visual fidelity.
3.2 Inverse Semantic Pyramid
@@
Preliminaries. Traditional convolutional neural networks (CNNs) naturally produce a pyramid of hierarchical bottom-up features $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H}{p\cdot 2^{l}}\times\frac{W}{p\cdot 2^{l}}\times C}\}$. Note that $l$ is the level index, $(H,W)$ is the image resolution, $C$ is the feature dimension, and $p$ is a down-sampling ratio ($e.g.$, $p=8$ in ResNet-50 [39]). In these pyramids, higher-resolution (lower-level) feature maps are rich in visual details and lower-resolution (higher-level) ones contain abstract semantic information. However, ViTs, splitting the image into coarse patches, only produce single-scale feature maps ($i.e.$, $\mathcal{F}^{0}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times C}$, where $e.g.,p=14$). Lacking such feature pyramids like CNNs hinders their performance on MLLM tasks requiring both fine-grained and high-level visual information. Therefore, how to construct a ViT-based feature pyramid with varying semantic granularity remains a problem.
@@
Visual detail injection module (VDIM). Due to the lack of high-resolution features, up-sampling ViT features becomes the necessary strategy to inversely construct the feature pyramid. Two simple approaches, (1) plain bilinear interpolation and (2) deconvolution network, can be adopted. By doubling and quadrupling the last-layer feature maps, a ViT-based feature pyramid $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times C},l=0,1,2\}$ is then constructed. While effective, directly up-sampling language-aligned features from CLIP-ViT hardly introduces precise visual details, resulting in suboptimal performance, illustrated in Table. 3. To address this, we design a VDIM to up-sample multi-modal semantic features guided by original image priors.
@@
Specifically, the objective of VDIM is to learn $(l-1)$ convolution layers on the image pyramid $\{\mathcal{I}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times 3}\}$ to capture high-frequency visual patterns of image texture for guiding the up-sampling process of semantic features, as shown in Fig. 3. For each input image, its $(l+1)$-th level features is defined as
@@
where $\mathrm{Up}(\cdot)$ denotes the up-sampling interpolation and $\mathrm{Conv}(\cdot)$ the convolutional operation on feature maps with customized kernel weights $\Theta^{l+1}$ learned on image $\mathcal{I}^{l+1}$.
Optimizing VDIM. We propose a multi-level reconstruction (MLR) loss between the higher-level feature maps $\{\mathcal{F}^{1},\mathcal{F}^{2}\}$ and the lowest one $\mathcal{F}^{0}$ as
$\mathcal{L}=\frac{1}{2}\sum_{l=1}^{2}\lVert\mathcal{F}^{0}-\mathrm{Down}(\mathcal{F}^{l};\Omega^{l})\rVert_{2}^{2},$ (2)
where $\mathrm{Down}(\cdot)$ is a down-sampling operation with trainable weights $\Omega^{l}$ in each level. The proposed MLR loss drives the feature pathway to capture low-level textures as well as to maintain multi-modal semantics during the fusion procedure.
Construction of inverse semantic pyramid (ISP). As shown in Fig. 3, VDIM acts as a progressive feature resolution expanding procedure conditioned on original image priors. During the inference, the resulting multi-level feature maps $\{\mathcal{F}^{0},\mathcal{F}^{1},\mathcal{F}^{2}\}$ form an ISP, which gathers a hierarchical multi-modal semantic representation with corresponding spatial resolutions supporting proper visual granularity.